what quanity x of a 75% acid solution must be mixed with a 25% solution to produce 500ml of a 40% solution?
Amount x=

Found 2 solutions by ptaylor, stanbon:
Answer by ptaylor(2198)

Let x=quantity of 75% solution needed
Then 500-x=quantity of 25% solution needed
Now we know that the amount of pure acid in the 75% solution (0.75x) plus the amount of pure acid in the 25% solution (0.25(500-x)) has to equal the amount of pure acid in the final solution (0.40*500). So our equation to solve is:
0.75x+0.25(500-x)=0.40*500 simplify
0.75x+125-0.25x=200 subtract 125 from each side
0.75x-0.25x=200-125 collect like terms
0.50x=75 divide each side by 0.50
x=150 ml----------------amount of 75% solution needed
CK
0.75*150+0.25*350=0.40*500
112.5+87.5=200
200=200
Hope this helps---ptaylor

Answer by stanbon(75887)

what quantity x of a 75% acid solution must be mixed with a 25% solution to produce 500ml of a 40% solution?
----
Quantity Equation::::: x + y = 500 ml
active ingredient Eq.:0.75x + 0.25y = 0.40*500 ml
---------------------------------------------------
Multiply thru 1st Equation by 25
Multiply thru the 2nd Eq. by 100
25x + 25y = 25*500
75x + 25y = 40*500
---------------------------
Subtract the 1st from the 2nd Eq.
50x = 15*500
x = 15*10
x = 150 ml (amount of 75% solution needed in the mixture)
Since x + y = 500, y = 350 ml (amount of 25% solution needed in the mixture)
==============================================================
Cheers,
Stan H.